OpenStudy (anonymous):

Bob bought some fruit candies weighing 5 grams each and some milk candies weighing 4 grams each.If he has 100 candies altogether and they weigh 480 grams in total, how many each kind does he have?

OpenStudy (shamim):

let no. of fruit candies is =x

OpenStudy (shamim):

so no. of milk candies r=100-x

OpenStudy (shamim):

weigh of fruit candies r=5x

OpenStudy (shamim):

weigh of milk candies=(100-x)4

OpenStudy (shamim):

so total weigh=5x+(100-x)4

OpenStudy (shamim):

so 5x+(100-x)4=480

OpenStudy (shamim):

now solve it for x